OpenStudy (anonymous):

Differentiate f(x) = ln((e^(4 x)(x - 1))/(2 x + 1) i got 4+(1)/(x-1)-(1)/(2x+1) is that right?

sam (.sam.):

\[\text{Use the chain rule, } \frac{d}{dx}\left(\log \left(\frac{e^{4 x} (x-1)}{2 x+1}\right)\right)=\frac{d\log (u)}{du} \frac{du}{dx} \text{, where } u=\frac{e^{4 x} (x-1)}{2 x+1} \text{ and } \frac{d\log (u)}{du}=\frac{1}{u} \]

sam (.sam.):

lol

OpenStudy (anonymous):

i cant see what you typed lol

sam (.sam.):

answer is \[\frac{e^{-4 x} \left(\left(4 e^{4 x} (x-1)+e^{4 x}\right) (2 x+1)-2 e^{4 x} (x-1)\right)}{(x-1) (2 x+1)}\] try to figure out? :)

OpenStudy (callisto):

Do you mean ln((e^(4 x)(x - 1)) / (2 x + 1) or ln ((e^(4 x)(x - 1))/(2 x + 1) ?

sam (.sam.):

Use chain and product rule

OpenStudy (anonymous):

i mean the first one

OpenStudy (anonymous):

ok i still get the same answer that i got

OpenStudy (callisto):

Is that your question (see the first line)

OpenStudy (anonymous):

yes that's the correct question

OpenStudy (callisto):

I made mistakes in the calculation...

OpenStudy (anonymous):

thats the right answer? if so, im way off..

OpenStudy (callisto):

Ask @.Sam.

sam (.sam.):

\[\frac{-8 x^2+4 x+1}{-2 x^2+x+1}\]

OpenStudy (anonymous):

thats the right answer?

sam (.sam.):

yeah, wait ill type the working

sam (.sam.):

\[\ln\frac{e^{4x}(x-1)}{2x+1}\] \[\ln(e^{4x}(x-1))-\ln(2x+1)\] differentiate \[\frac{1}{e^{4x}(x-1)}e^{4x}+(x-1)(4)(e^{4x})-\frac{1}{2x+1}(2)\] \[\frac{e^{4x}+(4x-4)(e^{4x})}{e^{4x}(x-1)}-\frac{2}{2x+1}\] \[\frac{e^{4x}(1+(4x-4)}{e^{4x}(x-1)}-\frac{2}{2x+1}\] \[\frac{4x-3}{x-1}-\frac{2}{2x+1}\] \[\frac{-8x^{2}+4x+1}{-2x^{2}+x+1}\]

OpenStudy (callisto):

 the question is 1 or 2?